OpenStudy (anonymous):

24t^3/15t^4*5t^8/3t^6

OpenStudy (anonymous):

\[24t ^{3}\times15t ^{4} \times5t ^{8}\div3t ^{6}\] i think that's right... thats the problem. \[so 24t ^{3} \times 15t ^{4} is 360 ^{7}\] because 24*15 is 360 and when you are multiplying exponents you add them. \[360t^{7} \times 5t^{8} is 1800t^{15}\] because again, you add exponents when you multiply them. \[1800t^{15}\div3t^{6}=600t^{9}\] because when you divide exponents, you subtract them. the answer is \[600t^{9}\] good luck on your test! trust me ive ben there. youll do fine.

OpenStudy (anonymous):

\[(24t^3/15t^4) * (5t^8/3t^6)\] \[(24t^3*5t^8)/(15t^4*3t^6)\] \[(120t^(3+8))/(45t^(4+6))\] \[120t^(11)/45t^(10)\]

OpenStudy (anonymous):

oh shoot i did that wrong. :( one seccc.

OpenStudy (anonymous):

the last part is: 120t^11/45t^10 = 120/45(t)^(11-10) = 120/45t^1 = 120/45t

OpenStudy (anonymous):

no no its ok i already got the answer

OpenStudy (anonymous):

the guy above me is right. If you are dividing exponents like this... \[5^{6}\div1^{5}\] you subtract the exponents and divide the base numbers. 5 to the first power! its the sam with multiplying. \[5^{6}\times 1^{5}\] yuo add the exponents to get 5 to the 11 power. i think that is what you were confused about. good luck on your test! trust me ive been there many times. youll do great!

OpenStudy (anonymous):

see i think u kinda go tthe question wrong but thnxs so much both of u gave u both a metal and fanned u guys

OpenStudy (anonymous):

thanks! i explained how to multiply and divide exponents above. good luck on your test!

OpenStudy (anonymous):

aww thnx so muck :)

OpenStudy (anonymous):

* much

OpenStudy (anonymous):

thanks cloverpetal, I often jump in like that also! kick myself later...!thanks for the medal maniqueen123

OpenStudy (anonymous):

np ;)\\